OpenStudy (anonymous):

Convert 3.4 radians into degrees

sam (.sam.):

\[X~radians \times \frac{180}{\pi}\] \[X ~degrees \times \frac{\pi}{180}\]

OpenStudy (anonymous):

well, which one is it? haha

OpenStudy (anonymous):

First one.